OpenStudy (anonymous):

a*.25 + (1-a)*.5 + 2* a * (1-a) *.25 *.5 *-.25

OpenStudy (yttrium):

What are we going to do here?

OpenStudy (anonymous):

set equal to 0 and solve for a

OpenStudy (yttrium):

Hmmm. Maybe I can show you it's simplified form and then you try to solve it. :) Here it is \[.25a+ .5(1-a) + 2*.25*.5a(1-a)-.25 = 0\] \[.25a+.5-.5a+.25a-.25a^2-.25 = 0\] Hence \[-.25a^2+.25 = 0\] Then you continue :p

OpenStudy (yttrium):

@perdue , do you get it?

OpenStudy (anonymous):

i was away for a second i'm going to try it now, thanks for simplifying it :)

OpenStudy (anonymous):

I'm getting that a = 1 but I know that a is supposed to = ,75

OpenStudy (yttrium):

Oopss. My answer is also 1 tho. Where did you get .75?

OpenStudy (anonymous):

i have the answer from the professor

OpenStudy (yttrium):

Maybe try to check your given again. I think there's something wrong with it. :)

OpenStudy (anonymous):

k thank you I'll take a look at it

OpenStudy (anonymous):

a*.25 + (1-a)*.5 + 2* a * (1-a) *.25 *.5 *(-.25) =0

OpenStudy (anonymous):

I'm getting something close to the .75 (.732), there could be some issues with rounding

OpenStudy (yttrium):

Is it something like this?

OpenStudy (anonymous):

.5 -.1875 -.0625a^2=0 is what I'm getting

OpenStudy (anonymous):

when you simplify the right side of the equation you get .0625

OpenStudy (anonymous):

multiplied by (a - a^2)

OpenStudy (anonymous):

.5 - .25a + .0625a -.0625a^2 =0

OpenStudy (anonymous):

I think it is a rounding error when you combine like terms and add the 'a coefficient terms' to both sides you get .5 = .1875a + .0625a^2

OpenStudy (yttrium):

isn't it be .25a+.5-.5a+.0625a-.0625a^2 = 0 ??

OpenStudy (anonymous):

yes that is right let me check it again, my algebra sucks

OpenStudy (anonymous):

but when you combine like terms - .25a -.5a + .0625a = (-.1875)

OpenStudy (anonymous):

so .5 -.1875a - .0625a^2 = 0

OpenStudy (yttrium):

How did you get your answer, did you used quadratic formula? Sorry huh, i do not have calculator here.

OpenStudy (anonymous):

actually I'm still not sure on the answer I'll let you know if/when I figure this out

OpenStudy (yttrium):

I have here the quadratic formula. Maybe this could help you. The quadratic equation in the form of ax^2 + bx + c = 0 may solve the roots thru this formula. \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] where a = -.0625a^2 b = -.1875a and c = .5 Just substitute the values over this formula

OpenStudy (anonymous):

thank you :)

OpenStudy (yttrium):

I hope it will lead to .75 haha lol